# literature.bib

### v2

@@ 4 Stability Analysis in Discrete Time
In continuous time, our theory currently provides us with a bound between the stationary distribution and the target $\mu^{\ast}$ via Theorem 3.11. In discrete time, we obtained bounds between the stationary distribution and the samples’ law. In order to make convergence statements between the samples and the target, we are currently working on a final coupling result to bridge the gap between discrete and continuous time stationary solutions. It will be reported in a future version of this manuscript.
@@ 5 Numerical Examples
In the following, we demonstrate the behaviour and our new results on ULPDA on several numerical examples, first small-scale ones to demonstrate the convergence behaviour in detail and subsequently larger imaging inverse problems. We always simulate (2.5) in a time interval $[0,T]$ by running the discretized algorithm (2.4) with small step sizes $\tau,\sigma$ for $T/\tau$ steps. The step sizes are coupled via $\lambda\tau=\sigma$ and are chosen small enough so that $\tau\sigma L^{2}\leq 1$, i.e. the assumptions of Theorem 4.1 are satisfied and the continuous-time dynamics are approximated well.
@@ 5.1 Revisiting the Quadratic Case
hence the target distribution is Gaussian $\mathrm{N}(0,c_{f}c_{g}(c_{g}k^{2}+c_{f})^{-1})$. For the numerical simulation, we choose $c_{f}=1,c_{g}=2,k=1.5$. We simulate the primal-dual Langevin dynamics (2.5) for different values of $\lambda$ by running ULPDA (2.4). We set $\theta=1$ and set the step sizes $\tau$ and $\sigma$ via the relations $\sigma/\tau=\lambda$ and $\sigma\tau k^{2}=c$ where $c\leq 1$ is required for stability. We set $c=10^{-4}$ in order to approximate closely the continuous time dynamics, resulting in step sizes $\tau=10^{-2}\sqrt{\lambda}k$ and $\sigma=10^{-2}\sqrt{\lambda^{-1}}k$. Figure 2 shows the joint distribution of primal and dual samples $(X^{n},Y^{n})$ produced by (2.4). As $\lambda$ grows, the joint distribution of primal and dual variable concentrates along the linear subspace $y=kx/c_{f}$ and the primal marginal distribution converges to the target.
Figure 2: Approximation of the target $\mu^{\ast}=\exp(-h)$ for a one-dimensional quadratic $h$ by (2.4) and its continuous time diffusion limit (2.5). From left to right, results for $\lambda=1$, $\lambda=10$ and $\lambda=100$ are displayed. The center scatter plot shows pairs $(X^{n},Y^{n})$, the background color indicates the mass of the concentrated joint target $\pi^{\ast}$ (zero is blue, the mass is concentrated along the white line $y=kx/c_{f}$). The marginal boxes show the densities of the primal and dual targets $\mu^{\ast}$ and $\nu^{\ast}=(\partial f\circ K)_{\#}\mu^{\ast}$ (dashed red), the marginals $\mu^{\text{stat}}_{\text{{cPD}}},\nu^{\text{stat}}_{\text{{cPD}}}$ of the invariant solution of primal-dual diffusion (black solid) and samples drawn using (2.4) (grey histograms). For increasing $\lambda$, the joint distribution of samples tends to concentrate and the primal marginal converges to the target, as proven in Theorem 3.11.
@@ 5.2.1 Toy Example in Two Dimensions
The distribution $\mu^{n}$ of iterate $X^{n}$ in ULPDA is approximated by running finitely many, here $10^{4}$, independent Markov chains in parallel and using the resulting empirical distribution $\hat{\mu}^{n}$. In Figure 3, we show scatter plots of the samples at time $T$, where $T$ is chosen large enough that the empirical measures $\hat{\mu}^{n}$ remain stationary both qualitatively and in Wasserstein distance to the target (see below). As comparison, we show samples drawn using the algorithm Prox-Sub proposed in [Habring2023]. Prox-Sub solves the optimality condition $y\in\partial f(Kx)$ exactly at every step (by assuming that the subgradient can be evaluated), and hence corresponds exactly to the concentrated limit $\lambda=\infty$ of ULPDA explored in Section 3.3. The authors of [Habring2023] showed the convergence of the distribution of Prox-Sub samples to the target in 2-Wasserstein distance, total variation distance and Kullback-Leibler divergence, where the bias decreases depends as usual on the step size of the algorithm.
@@
Figure 3: Scatter plots of $\hat{\mu}^{n}$ in the two-dimensional TV-denoising experiment. From left to right, the plots show approximations of the same posterior with values $\lambda=10^{1},10^{2},10^{3}$ and ‘$\lambda=\infty$’ (Prox-Sub). Horizontal and vertical axes correspond to the two coordinates in the primal variable $x\in\mathbb{R}^{2}$, the dual variable $y$ is not displayed. Grey dots show the samples at time $T$ produced by $10^{4}$ parallel chains, the black arrows indicate the samples’ mean and the principal components of their covariance (i.e. eigenvectors of the covariance matrix scaled in length by the corresponding eigenvalues). For reference, the blue arrows are the principal components of the posterior that should be approximated. The noisy observation $\tilde{x}$ is marked by a yellow triangle and the background color is a contour plot of the posterior density $\mu^{\ast}(x)$.
@@
Using numerical integration, in this low-dimensional example we are able to compute an estimate of the normalization constant $Z=\int_{\mathcal{X}}\exp(-f(Kx)-g(x))\,\mathrm{d}x$. Using this, we can access an approximation of the posterior density $\mu^{\ast}(x)$. This allows us to compute approximations of the Wasserstein distances of current iterates to the target $\mathcal{W}_{2}(\hat{\mu}^{n},\mu^{\ast})$ without the need to rely on a second sampler. We evaluate the distance at 20 equidistantly spaced points in $[0,T]$ (i.e. for $\hat{\mu}^{kN},k=1,\dots,20$ where $N=T/20\tau$). The distances are computed using the Python optimal transport module [Flamary2021] and are displayed in Figure 4. As expected, the samples drawn by the Prox-Sub algorithm are closest to the target distribution (their bias in Wasserstein distance is bounded by a constant multiple of the step size $\tau$). For finite $\lambda$, the distributions $\hat{\mu}^{n}$ exhibit an additional strong bias to the target due to the dualization, which becomes smaller as $\lambda$ grows.
@@
Figure 4: Wasserstein distances between the empirical measures $\hat{\mu}^{n}$ and the target $\mu^{\ast}$ in the two-dimensional TV-denoising experiment. We compare three different values of $\lambda$ and the algorithm Prox-Sub [Habring2023] corresponding formally to $\lambda=\infty$. The discretization via partial dualization of $f\circ K$ introduces a bias between samples and target that is significantly larger than the pure discretization bias of Prox-Sub. As $\lambda$ grows, the iteration approximates the purely primal setting and the bias is reduced.

### v3

@@ 4.2 Convergence to the Continuous Time Stationary Solution
The assumption of Lemma 4.6 that $\lVert A_{\text{{PD}}}(Z_{t})\rVert^{2}$ is bounded in expectation is a standard assumption on the drift in Langevin sampling algorithms. In the case of overdamped Langevin diffusion, it is satisfied if the drift coefficient, i.e. $\nabla h$ in (3.2), is assumed to be Lipschitz-continuous, see Lemma 2 of [Dalalyan2017a].
In the case of primal-dual diffusion, this condition puts growth requirements on $\nabla g$ and $\nabla f^{\ast}$ instead. If, e.g., there are constants $\alpha,\beta$ such that $\lVert\nabla g(x)\rVert\leq\alpha+\beta\lVert x\rVert$ and $\lVert\nabla f^{\ast}(y)\rVert\leq\alpha+\beta\lVert y\rVert$ for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, then the assumption is trivially satisfied. A further characterization of the stationary state $\pi_{\lambda}$ could help to replace this condition with more relaxed assumptions on $f$ and $g$.
The auxiliary inequality Lemma 4.5 and the bound on the discretization error in Lemma 4.6 allow us to give the following convergence result.
Theorem 4.8.
Let the assumptions of Lemma 4.6 be satisfied and assume further that $g,f^{\ast}$ are $\omega_{g}$- and $\omega_{f^{\ast}}$-strongly convex, respectively. Let the step sizes $\sigma,\tau$ and the extrapolation parameter $\theta$ be chosen such that $\theta\sigma\tau L^{2}\leq 1$ and
$\max\{(1+\omega_{g}\tau)^{-1},(1+\omega_{f^{\ast}}\sigma)^{-1}\}\leq\theta<1.$
Let the iterates $(X^{n},Y^{n})$ be generated by (1.4). Denoting $\mu^{n}$ the marginal distribution of $X^{n}$, we have
$W_{2}(\mu^{N},\mu_{\lambda})\leq\theta^{N/2}\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+C_{2}(\tau),$
and further, if $\theta\sigma\tau L^{2}<1$ for the distribution $\pi^{n}$ of the joint variable $(X^{n},Y^{n})$ it holds
$\mathcal{T}_{1,\bar{\lambda}}(\pi^{N},\pi_{\lambda})\leq\frac{\theta^{N/2}}{(1-\theta\sigma\tau L^{2})^{1/2}}\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+\frac{C_{2}(\tau)}{(1-\theta\sigma\tau L^{2})^{1/2}},$
with constants given by
$\bar{\lambda}\coloneqq\frac{1+\omega_{g}\tau}{1+\omega_{f^{\ast}}\sigma}\lambda,\quad C_{2}(\tau)\coloneqq\frac{(4dL_{\nabla g}^{2}\omega_{g}^{-1}+6dL^{2}(1+2\theta^{2})\omega_{f^{\ast}}^{-1})^{1/2}}{(1-\theta)^{1/2}}\tau+\mathcal{O}(\tau^{3/2}).$
Proof.
Let $(X^{n},Y^{n})$ be coupled with a process $\tilde{Z}_{t}=(\tilde{X}_{t},\tilde{Y}_{t})$, $t\geq 0$ by means of (4.6). Let $n\in\mathbb{N}_{0}$ be fixed and $t=n\tau$. Then
$\frac{2}{\tau}$ $\langle X^{n+1}-\tilde{X}_{t+\tau},\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\rangle$
$\leq\omega_{g}\lVert X^{n+1}-\tilde{X}_{t+\tau}\rVert^{2}+\frac{1}{\omega_{g}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2},$
as well as
$\frac{2}{\tau}$ $\langle Y^{n+1}-\tilde{Y}_{t+\tau},\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\rangle$
$\leq\omega_{f^{\ast}}\lVert Y^{n+1}-\tilde{Y}_{t+\tau}\rVert^{2}+\frac{1}{\omega_{f^{\ast}}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}.$
We now employ the auxiliary inequality Lemma 4.5. Inserting the quadratic lower bounds for the symmetric Bregman distances and the two bounds above, we obtain the inequality
$\lVert X^{n+1}-\tilde{X}_{t+\tau}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert X^{n+1}-\tilde{X}_{t+\tau}-X^{n}+\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}_{t+\tau}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}_{t+\tau}-X^{n}+\tilde{X}_{t}),Y^{n+1}-\tilde{Y}_{t+\tau}\rangle$
$\leq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle$
$\quad+\underbrace{\frac{1}{\omega_{g}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{p}_{s}-p^{\prime}+K^{T}(\tilde{Y}_{s}-\tilde{Y}_{t+\tau}))\,\mathrm{d}s\bigg{\rVert}^{2}}_{\eqqcolon R_{n}^{(1)}}+\underbrace{\frac{1}{\omega_{f^{\ast}}\tau^{2}}\bigg{\lVert}\int_{t}^{t+\tau}(\tilde{q}_{s}-\tilde{q}_{t+\tau}-K(\tilde{X}_{s}-\tilde{X}_{t,\theta}))\,\mathrm{d}s\bigg{\rVert}^{2}}_{\eqqcolon R_{n}^{(2)}},$
We now define
$\Delta^{n}$ $\coloneqq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}$
$\quad+2\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle.$
Given the assumptions on $\theta$ and the above inequality, it holds
$\Delta^{n+1}$ $\leq\lVert X^{n}-\tilde{X}_{t}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}_{t}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}_{t}-X^{n-1}+\tilde{X}_{t-\tau}),Y^{n}-\tilde{Y}_{t}\rangle+R_{n}^{(1)}+R_{n}^{(2)}$
$\leq\theta\Delta^{n}+R_{n}^{(1)}+R_{n}^{(2)}$
Iterating this inequality over $n=0,\dots,N-1$ we obtain
$\Delta^{N}\leq\theta^{N}\Delta^{0}+\sum_{n=0}^{N-1}\theta^{n}(R_{N-n-1}^{(1)}+R_{N-n-1}^{(2)})$
Denote $T=N\tau$. By lower bounding the remaining inner product in $\Delta^{N}$, this implies
$\lVert X^{N}-\tilde{X}_{T}\rVert_{\tau^{-1}+\omega_{g}}^{2}+(1-\theta\tau\sigma L^{2})\lVert Y^{N}-\tilde{Y}_{T}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}$
$\quad\leq\theta^{N}\left(\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}\right)+\sum_{n=0}^{N-1}\theta^{n}(R_{N-n-1}^{(1)}+R_{N-n-1}^{(2)})$
We now take expectations on both sides. The left side can be lower bounded by either $(\tau^{-1}+\omega_{g})\mathcal{W}_{2}^{2}(\mu^{N},\mu_{\lambda})$ or, in the joint variable, by $(\tau^{-1}+\omega_{g})(1-\theta\tau\sigma L^{2})\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{N},\pi_{\lambda})$. Since the initialization of the coupling was so far arbitrary (we only used that the marginal measures are $\pi_{0}$ and $\pi_{\lambda}$), we can choose the one that realizes the transport distance $\mathcal{T}_{1,\bar{\lambda}^{-1}}$ and obtain
$\mathbb{E}\left[\lVert X^{0}-\tilde{X}^{0}\rVert_{\tau^{-1}+\omega_{g}}^{2}+\lVert Y^{0}-\tilde{Y}^{0}\rVert_{\sigma^{-1}+\omega_{f^{\ast}}}^{2}\right]=(\tau^{-1}+\omega_{g})\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda}).$
Furthermore, by Lemma 4.6 we know that for all $n=0,\dots,N-1$ it holds
$\mathbb{E}\left[R_{N-n-1}^{(1)}+R_{N-n-1}^{(2)}\right]\leq\underbrace{(4L_{\nabla g}^{2}\omega_{g}^{-1}+6L^{2}(1+2\theta^{2})\omega_{f^{\ast}}^{-1})}_{\eqqcolon M}d\tau+\mathcal{O}(\tau^{2})$
This means we can use $\sum_{n=0}^{N-1}\theta^{n}\leq(1-\theta)^{-1}$ and after dividing by $\tau^{-1}+\omega_{g}$ we obtain the two estimates
$\mathcal{W}_{2}^{2}(\mu^{N},\mu_{\lambda})\leq\theta^{N}\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+\frac{Md}{1-\theta}\tau^{2}+\mathcal{O}(\tau^{3})$
and
$\mathcal{T}_{1,\bar{\lambda}^{-1}}^{2}(\pi^{N},\pi_{\lambda})\leq\frac{\theta^{N}}{1-\theta\sigma\tau L^{2}}\mathcal{T}^{2}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+\frac{Md}{(1-\theta\sigma\tau L^{2})(1-\theta)}\tau^{2}+\mathcal{O}(\tau^{3}),$
where we used that for $(\tau+\mathcal{O}(\tau^{2}))/(\tau^{-1}+\omega_{g})=\tau^{2}+\mathcal{O}(\tau^{3})$ for $\tau>0$. The statement of the theorem hence follows by taking the square root and using $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. ∎
In the last theorem, in the limit $N\to\infty$, the discrete time algorithm converges to its stationary distribution $\pi_{\lambda,\tau}$. From that, we immediately obtain the following bound between the invariant measures in discrete and continuous time.
Corollary 4.9.
Let the assumptions of Theorem 4.8 be satisfied. Then the following bounds hold:
$\mathcal{W}_{2}(\mu_{\lambda,\tau},\mu_{\lambda})$ $\leq C_{2}(\tau)+\mathcal{O}(\tau^{3/2}),$
$\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi_{\lambda,\tau},\pi_{\lambda})$ $\leq\frac{C_{2}(\tau)}{(1-\theta\sigma\tau L^{2})^{1/2}}+\mathcal{O}(\tau^{3/2}).$
The result Theorem 4.8 provides a bound between the samples produced by the algorithm and the corresponding stationary distribution in continuous time. While this result is typical in the Langevin algorithms literature, we here have to combine this with Theorem 3.11 in order to get a relation between samples and target distribution. The following result closes this final gap.
Corollary 4.10.
Let the assumptions of Theorems 3.11 and 4.8 be satisfied. Let $\pi^{n}$ be the distiribution of samples $(X^{n},Y^{n})$ generated by (1.4). Then the $x$-marginal $\mu^{n}$ of $\pi^{n}$ obeys
$\mathcal{W}_{2}(\mu^{N},\mu^{\ast})\leq\theta^{N/2}\mathcal{T}_{1,\bar{\lambda}^{-1}}(\pi^{0},\pi_{\lambda})+C_{1}(\lambda)+C_{2}(\tau),$
with the constants $C_{1},C_{2}$ from Theorems 3.11 and 4.8, respectively, which decay asymptotically like
$C_{1}(\lambda)$ $=D_{5}\lambda^{-1/2}+\mathcal{O}(\lambda^{-3/2}),\quad\textrm{as }\lambda\to\infty,$
$C_{2}(\tau)$ $=D_{6}\tau+\mathcal{O}(\tau^{3/2}),\quad\textrm{as }\tau\to 0.$
with constants gathered in $D_{5},D_{6}$. It immediately follows that $\mathcal{W}_{2}(\mu_{\lambda,\tau},\mu^{\ast})\leq C_{1}(\lambda)+C_{2}(\tau)$.
Corollary 4.11.
Let $(X^{n},Y^{n})$ be generated by the modified iteration (4.5), where the stochastic term $\xi$ is added in the primal variable before applying the proximal mapping. Under the same assumptions, the bounds of Theorem 4.8 and Corollaries 4.9 and 4.10 hold unchanged.
Proof.
As in the proof of Corollary 4.4, we have to modify the proof of the auxiliary inequality Lemma 4.5. It holds $p^{n+1}\in\partial g(X^{n+1})$, and accordingly we now choose $p^{\prime}\in\partial g(\tilde{X}_{t+\tau})$. Then, we start at the symmetric Bregman distance
$D_{g,\text{sym}}^{p^{n+1},p^{\prime}}(X^{n+1},\tilde{X}_{t+\tau})=\langle U^{n+1},p^{n+1}-p^{\prime}\rangle.$
The subsequent computations in the proof of Lemma 4.5 remain unchanged.
In order to bound the discretization error, the statement of Lemma 4.6 has to be changed to consider $p^{\prime}\in\partial g(X_{t+\tau})$ (instead of $p^{\prime}\in\partial g(X_{t+\tau}-\sqrt{2\tau}\xi)$). The error bound remains unchanged, the only difference in the proof appears in (4.9), where the term under the integral is now not $\mathbb{E}\left[\lVert X_{s}-X_{t+\tau}+\sqrt{2\tau}\xi\rVert^{2}\right]$, but $\mathbb{E}\left[\lVert X_{s}-X_{t+\tau}\rVert^{2}\right]$. Carrying out the subsequent computations in the same way, we arrive again at the upper bound $4L_{\nabla g}^{2}d\tau^{3}+\mathcal{O}(\tau^{4})$ for the primal discretization error. ∎
@@ 5 Numerical Examples
In the following, we demonstrate our results on the primal-dual Langevin sampling scheme on several numerical examples, first small-scale ones to demonstrate the convergence behaviour in detail and subsequently larger imaging inverse problems. We always simulate (2.6) in a time interval $[0,T]$ by running the discretized algorithm (1.4) with small step sizes $\tau,\sigma$ for $T/\tau$ steps. The step sizes are coupled via $\lambda\tau=\sigma$ and are chosen small enough so that $\tau\sigma L^{2}\leq 1$, i.e. the assumptions of Theorem 4.1 are satisfied and the continuous-time dynamics are approximated well.
@@ 5.1 Revisiting the Quadratic Case
hence the target distribution is Gaussian $\mathrm{N}(0,c_{f}c_{g}(c_{g}k^{2}+c_{f})^{-1})$. For the numerical simulation, we choose $c_{f}=1,c_{g}=2,k=1.5$. We simulate the primal-dual Langevin dynamics (2.5) for different values of $\lambda$ by running (1.4) (1.4). We set $\theta=1$ and set the step sizes $\tau$ and $\sigma$ via the relations $\sigma/\tau=\lambda$ and $\sigma\tau k^{2}=c$ where $c\leq 1$ is required for stability. We set $c=10^{-4}$ in order to approximate closely the continuous time dynamics, resulting in step sizes $\tau=10^{-2}\sqrt{\lambda}k$ and $\sigma=10^{-2}\sqrt{\lambda^{-1}}k$. Figure 3 shows the joint distribution of primal and dual samples $(X^{n},Y^{n})$ produced by (1.4). As $\lambda$ grows, the joint distribution of primal and dual variable concentrates along the linear subspace $y=kx/c_{f}$ and the primal marginal distribution converges to the target.
Figure 3: Approximation of the target $\mu^{\ast}=\exp(-h)$ for a one-dimensional quadratic $h$ by (1.4) and its continuous time diffusion limit (2.5). From left to right, results for $\lambda=1$, $\lambda=10$ and $\lambda=100$ are displayed. The center scatter plot shows pairs $(X^{n},Y^{n})$, the background color indicates the mass of the concentrated joint target $\pi^{\ast}$ (zero is blue, the mass is concentrated along the white line $y=kx/c_{f}$). The marginal boxes show the densities of the primal and dual targets $\mu^{\ast}$ and $\nu^{\ast}=(\partial f\circ K)_{\#}\mu^{\ast}$ (dashed red), the marginals $\mu^{\lambda},\nu^{\lambda}$ of the invariant solution of primal-dual diffusion (black solid) and samples drawn using (1.4) (grey histograms). For increasing $\lambda$, the joint distribution of samples tends to concentrate and the primal marginal converges to the target, as proven in Theorem 3.11.
@@ 5.2.1 Toy Example in Two Dimensions
The distribution $\mu_{n}$ of iterate $X^{n}$ in (1.4) is approximated by running finitely many, here $10^{4}$, independent Markov chains in parallel and using the resulting empirical distribution $\hat{\mu}_{n}$. In Figure 4, we show scatter plots of the samples at time $T$, where $T$ is chosen large enough that the empirical measures $\hat{\mu}_{n}$ remain stationary both qualitatively and in Wasserstein distance to the target (see below). As comparison, we show samples drawn using the algorithm Prox-Sub proposed in [Habring2023]. Prox-Sub solves the optimality condition $y\in\partial f(Kx)$ exactly at every step (by assuming that the subgradient can be evaluated), and hence corresponds exactly to the concentrated limit $\lambda=\infty$ of (1.4) explored in Section 3.3. The authors of [Habring2023] showed the convergence of the distribution of Prox-Sub samples to the target in 2-Wasserstein distance, total variation distance and Kullback-Leibler divergence, where the bias decreases depends as usual on the step size of the algorithm.
@@
Figure 4: Scatter plots of $\hat{\mu}_{n}$ in the two-dimensional TV-denoising experiment. From left to right, the plots show approximations of the same posterior with values $\lambda=10^{1},10^{2},10^{3}$ and ‘$\lambda=\infty$’ (Prox-Sub). Horizontal and vertical axes correspond to the two coordinates in the primal variable $x\in\mathbb{R}^{2}$, the dual variable $y$ is not displayed. Grey dots show the samples at time $T$ produced by $10^{4}$ parallel chains, the black arrows indicate the samples’ mean and the principal components of their covariance (i.e. eigenvectors of the covariance matrix scaled in length by the corresponding eigenvalues). For reference, the blue arrows are the principal components of the posterior that should be approximated. The noisy observation $\tilde{x}$ is marked by a yellow triangle and the background color is a contour plot of the posterior density $\mu^{\ast}(x)$.
@@
Using numerical integration, in this low-dimensional example we are able to compute an estimate of the normalization constant $Z=\int_{\mathcal{X}}\exp(-f(Kx)-g(x))\,\mathrm{d}x$. Using this, we can access an approximation of the posterior density $\mu^{\ast}(x)$. This allows us to compute approximations of the Wasserstein distances of current iterates to the target $\mathcal{W}_{2}(\hat{\mu}_{n},\mu^{\ast})$ without the need to rely on a second sampler. We evaluate the distance at 20 equidistantly spaced points in $[0,T]$ (i.e. for $\hat{\mu}_{kN},k=1,\dots,20$ where $N=T/20\tau$). The distances are computed using the Python optimal transport module [Flamary2021] and are displayed in Figure 5. As expected, the samples drawn by the Prox-Sub algorithm are closest to the target distribution (their bias in Wasserstein distance is bounded by a constant multiple of the step size $\tau$). For finite $\lambda$, the distributions $\hat{\mu}_{n}$ exhibit an additional strong bias to the target due to the dualization, which becomes smaller as $\lambda$ grows.
@@
Figure 5: Wasserstein distances between the empirical measures $\hat{\mu}_{n}$ and the target $\mu^{\ast}$ in the two-dimensional TV-denoising experiment. We compare three different values of $\lambda$ and the algorithm Prox-Sub [Habring2023] corresponding formally to $\lambda=\infty$. The discretization via partial dualization of $f\circ K$ introduces a bias between samples and target that is significantly larger than the pure discretization bias of Prox-Sub. As $\lambda$ grows, the iteration approximates the purely primal setting and the bias is reduced.